OpenStudy (anonymous):

a+9b=20 a+3b=2 how would you use elimination method for this

hero (hero):

Hint: Multiply the second equation by -3

OpenStudy (anonymous):

a+9b=20 a+3b=2 try to eliminate a , we have 6b=18 so b is ?

OpenStudy (anonymous):

You could also multiply one of the equation by -1 and then solve for b.

OpenStudy (anonymous):

I am confused

OpenStudy (anonymous):

a+9b=20 a+3b=2 (a+9b=20)*(-1) a+3b=2 -a-9b=-20 a+3b=2 Add the two equations together and you'll get -6b=-18 b=(-18/-6) b=3 Plug b=3 back into one of the equations. a+3b=2 a+3(3)=2 Solve for a a = 2-3(3) a = 2-9 a = -7

OpenStudy (anonymous):